OpenStudy (anonymous):

Determine which of the following statements is true concerning the values described in column #1 and column #2. Column #1 Column #2 The x-coordinate of the vertex of the graph of y = −2x2 − 4x + 12 The x-coordinate of the vertex of the graph of y = x2 − 4x + 3 The value found in column #1 is greater than the value found in column #2. The value found in column #1 is less than the value found in column #2. The value found in column #1 is equivalent to the value found in column #2. The relationship between column #1 and column #2 cannot be determined by the

OpenStudy (e.cociuba):

y = -2x^2 - 4x + 12 = -2(x^2 + 2x) + 12 = -2(x^2 + 2x + 1) + 12 + 2 = -2(x + 1)^2 + 14 y = x^2 - 4x + 3 = (x^2 - 4x) + 3 = (x^2 - 4x + 4) + 3 - 4 = (x - 2)^2 - 1 The vertex is (-1, 14) in the first; (2, -1) in the second. ANSWER: The value found in column #1 is less than the value found in column #2.

OpenStudy (anonymous):

thx @e.cociuba

OpenStudy (dancergirl45):

was this right